answer to
Through (-3,4);horizontal is it 3x+4y=0?
" " (-2,8) m = -1 is it 2x + y= -8?
" " (-4,7); parallell to 3x+4y = 4
" " (-4,7); perpendicular to x-2y = 0
" " (2,-7) and (3,-1) is it x-6y=44?

Thank You, really struggling through this stuff!!!

Answer by Alan3354(69443)

answer to
Through (-3,4);horizontal is it 3x+4y=0?
Since it's horizontal (parallel to the x-axis), y does not change.
That makes it y = 4.
-----------------------
" " (-2,8) m = -1 is it 2x + y= -8?
For these, always use
y-y1 = m*(x-x1) where (x1,y1) is the given point.
y-8 = -1*(x+2)
y-8 = -x-2
x+y = 6
-------------
" " (-4,7); parallell to 3x+4y = 4
In this one, you have to find the slope of 3x+4y=4. Put it in slope-intercept form (that means solve for y).
3x+4y = 4
4y = -3x+4
y = (-3/4)x + 1
m = -3/4
----------
Then it's the same as the one above.
y-y1 = m*(x-x1)
------------------
" " (-4,7); perpendicular to x-2y = 0
Find the slope of x-2y=0
y = (1/2)x --> m = 1/2
The slope of lines perpendicular will be the negative inverse.
m = -2
----------
Then it's the same as above.
-----------------
" " (2,-7) and (3,-1) is it x-6y=44?
Find the slope.
m = diffy/diffx = (-1+7)/(3-2)
m = 6
Then, use either point and same as above.
y+1 = 6*(x-3)
y+1 = 6x-18
6x-y = 17